### 18 Dec 2023 v1

# C.V. Raman Avenue, Bengaluru 560012, India

Abstract
Building upon recent research in spin systems with non-local interactions, this study investigates operator growth using the Krylov complexity in different non-local versions of the Ising model. We find that the non-locality results in a faster scrambling of the operator to all sites. While the saturation value of Krylov complexity of local integrable and local chaotic theories differ by a significant margin, this difference is much suppressed when non-local terms are introduced in both regimes. This results from the faster scrambling of information in the presence of non-locality. In addition, we investigate the behavior of level statistics and spectral form factor as probes of quantum chaos to study the integrability breaking due to non-local interactions. Our numerics indicate that even in the non-local case, Krylov complexity can distinguish between different underlying theories as well as different degrees of non-locality.
1 Introduction
The study of operator growth using Krylov complexity (also referred to as ‘K-complexity’) has emerged as a powerful framework for investigating the behavior of quantum systems and their underlying dynamics PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a ; rabinoviciOperatorComplexityJourney2021b . K-complexity is a measure of the delocalization of a local initial operator evolving under Heisenberg evolution with respect to the Hamiltonian PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a ; rabinoviciOperatorComplexityJourney2021b ; PhysRevB.102.085137 ; jianComplexityGrowthOperators2021 . It is conjectured to grow at most exponentially generically non-integrable systems PhysRevX.9.041017 . This exponential growth of K-complexity can be used to extract the Lyapunov exponent PhysRevX.9.041017 ; barbonEvolutionOperatorComplexity2019a , establishing a connection with out-of-time-ordered-correlators (OTOC) maldacenaBoundChaos2016 ; PhysRevResearch.2.043234 . Further studies have shown a relation between K-complexity and chaos in context of various models such as Ising models rabinovici_krylov_2022 ; rabinoviciKrylovLocalizationSuppression2022 , Sachdev-Ye-Kitaev (SYK) models jianComplexityGrowthOperators2021 ; bhattacharjeeKrylovComplexityLarge2023 ; heQuantumChaosScrambling2022 , quantum field theories caputaOperatorGrowth2d2021 ; khetrapalChaosOperatorGrowth2023 ; kunduStateDependenceKrylov2023 ; camargoKrylovComplexityFree2023 ; avdoshkinKrylovComplexityQuantum2022 ; Erdmenger:2023wjg , many-body localization system 10.21468/SciPostPhys.13.2.037 , and open quantum systems bhattacharyaOperatorGrowthKrylov2022b ; bhattacharjeeOperatorGrowthOpen2023 ; bhattacharyaKrylovComplexityOpen2023d ; PhysRevResearch.5.033085 .
Among different systems, of particular interest are the so-called fast scramblers for which scrambling time is logarithmic in the number of degrees of freedom. The examples include black holes Yasuhiro_Sekino_2008 ; lashkariFastScramblingConjecture2013 ; maldacenaBoundChaos2016 , conjectured to be the fastest scramblers in nature, SYK PhysRevLett.70.3339 and other related holographic models. The previous studies PhysRevA.94.040302 ; PhysRevA.99.051803 ; PhysRevLett.123.130601 ; belyansky2020minimal suggest that local chaotic dynamics along with non-local interactions are sufficient to give rise to the phenomena termed as “fast scrambling”. These developments have motivated the study of quantum information scrambling in non-local systems PhysRevResearch.2.043399 ; wanisch2022quantum as well as experimental proposal for probing fast scramblers with simpler models PhysRevLett.123.130601 ; belyansky2020minimal ; PhysRevResearch.2.043399 ; PhysRevB.107.L081103 .
In this study, we characterize the growth of K-complexity in non-local systems. It has been established that K-complexity can distinguish between the integrable and chaotic regimes in local systems. To be precise, the complexity growth rate and saturation value are known to be significantly higher for chaotic systems as compared to the integrable ones rabinoviciOperatorComplexityJourney2021b ; rabinovici_krylov_2022 ; rabinoviciKrylovLocalizationSuppression2022 . Here, we ask the question of how these characteristics get modified once non-local interactions are turned on within these regimes. While chaotic systems with non-local interactions can exhibit fast scrambling properties, the non-local terms in the otherwise integrable cases are also expected to accelerate the operator growth and therefore begs some investigation as to whether the notions of local integrability and chaos remain distinguishable in the presence of non-locality.
Apart from K-Complexity, we also study the two widely used measures to probe chaos: Level Statistics (Appendix A.1) and Spectral Form Factor (Appendix A.2). The symmetry-reduced level statistics is known to exhibit Poisson distribution for integrable regimes, while in chaotic regimes, it shows the Wigner-Dyson distribution 10.1007/978-3-0348-8266-8_36 ; PhysRevLett.52.1 ; PhysRevLett.110.084101 . On the other hand, the spectral form factor (SFF) is found to exhibit an explicit dip-ramp-plateau behaviour for chaotic evolution in contrast to the integrable regime where higher amount of fluctuations wash out the dip-ramp-plateau behaviour cotlerBlackHolesRandom2017 ; PhysRevD.95.126008 ; winerSpectralFormFactor2022 . [^2] This might vary from model to model. In some cases, an effective ramp appears, but it is not as evidently clear as it is for chaotic ones. We study these quantities to compare their behaviour in the presence of non-local terms in the Hamiltonian and analyse how the introduction of non-locality changes their nature.
Our results indicate that with increasing degrees of non-locality, the integrable evolution becomes similar to chaotic ones. We study the three diagnostics for characterizing chaos for various models which indicates that non locality apparently makes all regimes chaotic. However, a closer look at the plots reveals that non-local terms with local integrable operator evolution can still be distinguished from non-local terms with local chaotic operator evolution using K-complexity. Moreover, we also observe that the increase in non-locality is captured by the initial growth rate of the K-complexity. To have a clearer understanding of how non-locality changes behavior of Krylov complexity, we also study “mixed field” spin chains wanisch2022quantum , where we vary the degree of non-locality and study how it increases the slope of the Lanczos ascent.
The rest of the paper is structured as follows. In section 2, we briefly review the notion of K-complexity, and it’s characteristics for local Hamiltonians, discussing both integrable and non-integrable regimes. In section 3, we introduce several non-local models and discuss distinguished features present in the K-complexity profiles. Section 3.1 consists of the numerical findings for non-locality introduced to an otherwise local integrable and chaotic version of the transverse field Ising model. In sections 3.2 and 3.3, we present the models and numerics when we deal with varying degrees of non-locality by tuning a parameter in transverse field Ising model and XXZ spin chain respectively. Finally, we conclude in section 4 discussing the implications of our results. In Appendix we study the remaining probes, namely the level statistics and the spectral form Factor for both local and non-local models to be able to compare their behavior with Krylov complexity.
2 Review for local Hamiltonians
In this section, we briefly review Krylov complexity and the features of operator growth for local Hamiltonians. As part of our exploration, we will focus on how this behaves within the context of the transverse-field Ising model, considering both integrable and non-integrable scenarios. For our analysis, we will take the local Hamiltonian to be a one-dimensional transverse-field Ising model with open boundary conditions, which has the form,
$H_{\mathrm{TFIM}}=-\sum_{j=1}^{N-1}\sigma^{z}_{j}\sigma^{z}_{j+1}-g\sum_{j=1}^{N}\sigma^{x}_{j}-h\sum_{j=1}^{N}\sigma^{z}_{j}\,,$ (1)
where $g$ and $h$ are the coupling parameters. When $h=0$, the Hamiltonian remains integrable for all values of $g$, as it can be mapped to the free-fermionic model sachdev_2011 . However, when both $g$ and $h$ are nonzero, the system departs from integrability. To investigate the integrable regime, we set $g=1$ and $h=0$, while for the chaotic regime, we choose $g=-1.05$ and $h=0.5$ PhysRevLett.106.050405 . The model has a parity symmetry for all values of coupling parameters, and additional $Z$-reflection symmetry when $h=0$.
2.1 Operator dynamics in Krylov basis
We begin with a brief review of the operator dynamics in Krylov spacePhysRevX.9.041017 ; rabinoviciOperatorComplexityJourney2021b ; barbonEvolutionOperatorComplexity2019a . Consider the evolution of an operator with seed $\mathcal{O}_{0}$ at $t=0$ under the time-independent Hamiltonian $H$. At any time $t$, the time-evolved operator $\mathcal{O}(t)$ under Heisenberg evolution can be written as
$\mathcal{O}(t)=e^{iHt}\mathcal{O}_{0}e^{-iHt}=e^{i\mathcal{L}t}\mathcal{O}_{0}=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}_{0}$ (2)
where $\mathcal{L}$ is Liouvillian superoperator, which satisfies the relation $\mathcal{L}\mathcal{O}=[H,\mathcal{O}]$. This motivates the definition of Krylov space associated with the operator $\mathcal{O}$ as the minimal subspace of operator space that contains the time evolution of $\mathcal{O}$ at all times. Therefore, the Krylov space is obtained from repeated action of Liouvillian $\mathcal{L}^{n}\mathcal{O}$ :
$\mathcal{H}_{\mathcal{O}}=\text{span}\{\mathcal{L}^{n}\mathcal{O}\}^{\infty}_{n=0}=\text{span}\{\mathcal{O},[H,\mathcal{O}],[H,[H,\mathcal{O}]],\ldots\}$ (3)
Once the Krylov space is obtained, an orthonormal basis is then formed using a choice of inner product on the operator space. This is achieved using the Lanczos algorithm, an instance of the Gram-Schmidt process PhysRevX.9.041017 ; baiTemplatesSolutionAlgebraic1987 . For our purpose, we will use the infinite-temperature inner product, also known as Frobenius inner product :
$(\mathcal{A}|\mathcal{B})=\frac{1}{D}\text{Tr}\left[\mathcal{A}^{\dagger}\mathcal{B}\right],\ \ \ \lVert\mathcal{A}\rVert=\sqrt{(\mathcal{A}|\mathcal{A})},$ (4)
where $D=\text{Tr}[\textbf{I}]$ is the trace of the identity matrix of appropriate dimension. In the orthonormal basis obtained from Lanczos iteration, the Liouvillian takes a tridiagonal form :
$\mathcal{L}=\begin{pmatrix}0&b_{1}&&&&0\\ b_{1}&0&b_{2}&&&\\ &b_{2}&0&\ddots&&\\ &&\ddots&\ddots&b_{K-2}&\\ &&&b_{K-2}&0&b_{K-1}\\ 0&&&&b_{K-1}&0\\ \end{pmatrix}$ (5)
where K is dimension of Krylov space, constrained to $11\leq K\leq D^{2}-D+$ 1 rabinoviciOperatorComplexityJourney2021b . Eq. (5) can be recast into the following form,
$\mathcal{L}|\mathcal{O}_{n})=b_{n}|\mathcal{O}_{n-1})+b_{n+1}|\mathcal{O}_{n+1})$ (6)
We can expand the time-evolving operator in the Krylov basis as :
$|\mathcal{O}(t))=\sum_{n=0}^{K-1}i^{n}\phi_{n}(t)|\mathcal{O}_{n})$ (7)
where $\phi_{n}(t)$ are time-dependent probability amplitudes associated with the Krylov chain. The wavefunctions satisfy the recursive differential equation, followed by the Heisenberg equation, which takes the form
$\dot{\phi}_{n}(t)=b_{n}\phi_{n-1}(t)-b_{n+1}\phi_{n+1}(t)$ (8)
where $\phi_{-1}(t)=0$ and $\phi_{n}(0)=\delta_{n0}$. The unitarity of operator evolution implies a normalization condition on the probability amplitudes,
$\sum_{n=0}^{K-1}|\phi_{n}(t)|^{2}=1$ (9)
In Eq. (8), we can think of the Lanczos coefficients $b_{n}$ as the hopping amplitudes for the seed $O_{0}$ to traverse in the Krylov Space and $\phi_{n}$ as the probability amplitude associated with site $n$. The average position of the distribution on the Krylov chain is called K-complexity and is given by
$C_{K}(t)=\sum_{n=0}^{K-1}n|\phi_{n}(t)|^{2}$ (10)
We can also define the complexity operator hornedal_ultimate_2022 ,
$\mathcal{K}=\sum_{n=0}^{K-1}n|\mathcal{O}_{n})(\mathcal{O}_{n}|$ (11)
which plays the role of the position operator in the Krylov chain so that the Krylov complexity can be written as the expectation value of the complexity operator $\mathcal{K}$
$C_{K}(t)=(\mathcal{O}(t)|\mathcal{K}|\mathcal{O}(t))\,.$ (12)
Figure 1: Left: The Lanczos-sequence for the TFIM Hamiltonian in Eq. (1) computed for $L=7$ spins with open boundary conditions for fixed parity $P=+1$ in (a) Integrable (b) Chaotic limits. Right: Krylov complexity for TFIM Hamiltonian (1) computed for $L=7$ spins with open boundary conditions for fixed parity $P=+1$ in Integrable and Chaotic limits. Outset: Full-time range computed. Inset: Zoom in at early times.
To study K-complexity we will use open boundary conditions and focus on a local operator $\mathcal{O}$ which respects the parity symmetry and keeps the computation within the chosen sector,
$\mathcal{O}=S^{z}_{i}+S^{z}_{N-i+1},$ (13)
where $i$ is chosen somewhere close to the center of the chain. Fig. 1 shows the Lanczos sequence $b_{n}$ and K-complexity for TFIM Hamiltonian (1) computed for $L=7$ for fixed parity $P=+1$ sector for the operator $\mathcal{O}=S^{z}_{4}$. The study of Lanczos sequence and Krylov complexity in integrable and chaotic limits have extensively been done in previous studies for different systems rabinovici_krylov_2022 ; iizukaKrylovComplexityIP2023 ; bhattacharyyaOperatorGrowthKrylov2023 ; camargoSpectralKrylovComplexity2023 . The Lanczos coefficient features sublinear growth in the integrable limit while linear growth in the chaotic limit, followed by saturation, and the descent, while K-complexity transitions from exponential growth at very early time to linear increase followed by saturation rabinoviciOperatorComplexityJourney2021b ; rabinovici_krylov_2022 . The saturation value of K-complexity is significantly large for local chaotic evolution as compared to the local integrable one. The suppressed complexity saturation of integrable cases results from higher fluctuations in Lanczos coefficients during the descent phase. This fluctuating descent period is dubbed as the Krylov localization.
3 Non-local models and numerical findings:
In this section, we will study operator growth in non-local systems. We compare the results with the local counterpart to demonstrate the effect of non-locality on these quantities. For these Hamiltonians, it is, in general, expected that due to non-local terms, the Liouvillian will also be non-local. Now, since the notion of integrability is believed to be closely related to locality PhysRevResearch.2.043399 ; wanisch2022quantum , the evolution of the operator under the non-local Liouvillian is expected to show chaotic behavior even in the case when the local part of the Hamiltonian is integrable. Therefore it is expected that the non-local terms in the Hamiltonian will result in the loss of its integrability.
We study two types of non-local terms in this paper: i) All-to-all couplings with the same interaction strength which we call the non-local transverse field Ising model, and ii) All-to-all couplings with varying interaction strength, where the coupling between two sites decay as a power of the distance between the sites. We call this the transverse mixed field Ising model and by tuning the power, we vary the degree of non-locality in this model.
3.1 Non-local Transverse Field Ising Model
We consider the fast scrambling spin-1/2 Hamiltonian introduced in belyansky2020minimal ,
$\mathcal{H}=\mathcal{H}_{\text{local}}-\frac{\gamma}{\sqrt{N}}\sum_{i<j}\sigma^{z}_{i}\sigma^{z}_{j},$ (14)
where $\sigma^{z}_{i}$ is the Pauli $z$ operator acting on site $i$ and $\mathcal{H}_{\text{local}}$ is Hamiltonian with only local interactions. In our study, we will consider the local Hamiltonian as a one-dimensional transverse-field Ising model with open boundary conditions introduced in Sec. 2. At this point, it is important to emphasize the terminology used throughout this literature. We will later observe from the results of operator growth as well as level-statistics and spectral form factor (discussed in the appendices A.1 and A.2), that non-locality can break the integrability of the system. While referring to a system whose local part of the Hamiltonian in Eq. (14) is integrable, we will call it non-local integrable Hamiltonian even though non-locality breaks integrability, and the model is no longer integrable.
Figure 2: Comparison between Lanczos-coefficients for local and non-local TFIM Hamiltonian with non-local parameter $\gamma=0.5$, computed for $N=7$ spins in the $P=+1$ sector for Integrable and Chaotic limits of local Hamiltonian. Top row: complete spectrum of Lanczos coefficients, Bottom row: Initial lanczos coefficients.
Figure 3: Krylov complexity for TFIM Hamiltonian (1) computed for $L=7$ spins with open boundary conditions for fixed parity $P=+1$, and non-locality parameter $\gamma=0.5$ in Left: Integrable Right: Chaotic limits. Top: Full-time range computed. Bottom: Zoom in at early times.
We studied the non-local TFIM Hamiltonian with $L=7$ sites and non-local parameter $\gamma=0.5$. The numerical results for Lanczos-sequence and Krylov complexity are shown in Fig. 2 and 3. The global picture emerging from these results can be summarized in the following points:
  • Lanczos sequence The initial growth of Lanczos coefficients is faster for the non-local cases as compared to local cases (in both integrable and chaotic limits). The peak value of the Lanczos sequence is more than that of the corresponding local case. The overall Lanczos profile for the non-local integrable case resembles the form of the local chaotic case. The fluctuations in the decaying part of the Lanczos sequence are also small for the non-local integrable case which results in a high saturation value of K-complexity.
  • K-complexity This shows a transition from exponential growth to linear increase starting from the scrambling time. This transition time is lower for non-local models in both integrable and chaotic regimes. Furthermore, the initial growth rate is also greater in non-local cases compared to their local counterparts. At exponentially late times the complexity saturates at half of the Krylov dimension for the chaotic cases local and non-local cases, since by then the operator is uniformly distributed over the Krylov basis. It is also worth noting that even after the introduction of non-local terms to the otherwise local integrable Hamiltonian, the complexity saturation value does not reach exactly the value of the local chaotic Hamiltonian. The saturation value for non-local terms $+$ local integrable Hamiltonian is close to $1500$ (much higher compared to local integrable which was $500$) whereas the the saturation value for non-local terms $+$ local chaotic Hamiltonian is close to $2000$ (almost similar to the local chaotic case). The late time saturation value of the chaotic Hamiltonian is the same for both local and non-local cases, although the non-local chaotic Hamiltonian has a large initial growth compared to the local chaotic Hamiltonian.
3.2 Transverse Mixed Field Ising Model
Consider the one-dimensional transverse mixed field Ising model of $N$ spins with open boundary conditions
$\mathcal{H}_{\alpha}=-\sum_{i<j}J^{\alpha}_{ij}\sigma^{z}_{i}\sigma^{z}_{j}-g\sum_{i}\sigma^{x}_{j}-h\sum_{j}\sigma^{z}_{j}$ (15)
where $J^{\alpha}_{ij}$ is interaction strength between spins at position $i$ and $j$ which assume to follow powerlaw $1/\kappa\cdot J/|i-j|^{\alpha}$. Throughout our study, we take $J=1$, and $\kappa=1$. In previous studies wanisch2022quantum , the information scrambling, using out-of-time-order correlators (OTOCs) as a probe, has been studied in such a model with a variation of non-local parameter $\alpha$. It’s shown that for $\alpha>2$, the dynamics effectively resemble local dynamics, while for smaller exponents, the dynamics become non-local. Here, we will confine ourselves to studying the initial Lanczos sequence for the model with varying exponent $\alpha$.
Figure 4: The Lanczos-sequence for the Transverse Mixed Field Ising Model Hamiltonian in Eq. (15) computed for $L=13$ spins with open boundary conditions for fixed parity $P=+1$ in (a) Integrable (b) Chaotic limits with varying non-local exponent $\alpha$ values.
Figure 5: Krylov complexity for the Transverse Mixed Field Ising Model Hamiltonian in Eq. (15) computed for $L=13$ spins with open boundary conditions for fixed parity $P=+1$ in (a) Integrable (b) Chaotic limits with varying non-local exponent $\alpha$ values.
In Fig. 4, we showed the initial Lanczos coefficients $b_{n}$ for $L=13$ with varying exponent $\alpha\in[0.1,0.5,1.0,1.5,2.0,2.5]$. The initial operator is chosen to be $S^{z}_{7}$. In both limits integrable and chaotic, we find the increasing slope and saturation value in initial growth with the decrease in exponent $\alpha$ and therefore increase in non-locality. This ensures that the initial (pre-scrambling) growth rate of the Krylov complexity always increases in the presence of non-locality (See Fig. 5). In Figure 6, we showed the linear growth rate calculated using linear fit ($b_{n}=\delta n+c$) in initial Lanczos coefficients.
Figure 6: The linear growth rate $\delta$ plotted for different non-local exponent $\alpha$ values for Lanczos coefficients in Fig. 4.
3.3 Mixed field XXZ Hamiltonian
In this section, we will introduce mixed-field XXZ Hamiltonian analogous to mixed-field TFIM. The XXZ Hamiltonian contains nearest-neighbour interaction terms which are used to describe the behavior of a system of interacting spin-1/2 particles in a magnetic field and the Hamiltonian can be written as,
$H_{\text{XXZ}}=\frac{1}{4}\sum_{i=1}^{N-1}J\,(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y})+\frac{J_{zz}}{4}\,\sigma_{i}^{z}\sigma_{i+1}^{z}\,.$ (16)
The XXZ Hamiltonian commutes with total spin operator $M$ in the $z$-direction and is also invariant under reflection with respect to the edge of the chain i.e. under parity operator $P$. In previous studies Santos_2004 ; PhysRevE.84.016206 ; PhysRevB.80.125118 ; PhysRevB.98.235128 ; PhysRevB.102.075127 ; PhysRevX.10.041017 ; gubin_quantum_2012 ; rabinovici_krylov_2022 , it was shown that the integrability can be broken by addition local term
$\begin{split}H_{d}&=S_{j}^{z}\,.\end{split}$ (17)
The mixed field XXZ is an extension of the XXZ Hamiltonian where we add varying non-local strength analogous to the case of mixed field TFIM.
$\mathcal{H}^{(\alpha)}_{XXZ}=\frac{1}{4}\sum_{i=1}^{N-1}J^{\alpha}_{ij}\,(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y})+\frac{\tilde{J}^{\alpha}_{ij}}{4}\,\sigma_{i}^{z}\sigma_{i+1}^{z}$ (18)
where $J^{\alpha}_{ij}$ and $\tilde{J}^{\alpha}_{ij}$ are interaction strength between spins at position $i$ and $j$. These follow powerlaw behavior, $1/\kappa\cdot J/|i-j|^{\alpha}$, and $1/\kappa\cdot J_{zz}/|i-j|^{\alpha}$ respectively. Note that this model reduces to local XXZ model in the limit $\alpha\rightarrow\infty$. We consider the following non-local interpolating Hamiltonian:
$\mathcal{H}_{\text{non-local}}=\mathcal{H}^{(\alpha)}_{XXZ}+\epsilon_{d}H_{d}\,.$ (19)
In our study, we will consider $H_{d}=S^{z}_{(N+1)/2}$, and therefore, both the symmetry-breaking terms keep the symmetries intact. In further calculations, we fix $J=1$, and $J_{zz}=1.1$ for all cases.
Figure 7: The Lanczos-sequence for the Transverse Mixed Field XXZ Model Hamiltonian in Eq. (15) computed for $L=12$ spins with open boundary conditions for fixed parity $P=+1$ and $M=5$ in (a) $\epsilon_{d}=0$ (b) $\epsilon_{d}=0.5$ limits with varying non-local exponent $\alpha$ values. The initial operator is chosen to be $S^{z}_{6}+S^{z}_{7}$.
Figure 8: Krylov complexity for the Transverse Mixed Field XXZ Model Hamiltonian in Eq. (15) computed for $L=12$ spins with open boundary conditions for fixed parity $P=+1$ and $M=5$ in (a) $\epsilon_{d}=0$ (b) $\epsilon_{d}=0.5$ limits with varying non-local exponent $\alpha$ values.
We study the initial Lanczos sequence for $L=12$ with varying exponent $\alpha$ for zero and non-zero value of $\epsilon_{d}$ (See Fig. 7). The initial operator $\mathcal{O}$ is chosen such that it respects the two symmetries of the model, i.e., parity and total spin in $z$-direction
$\mathcal{O}=S^{z}_{i}+S^{z}_{N-i+1}$ (20)
where $i$ is chosen to be near the center of the chain. We find the increasing slope and saturation value in initial growth with the decrease in exponent $\alpha$ and, therefore increase in non-locality. These results are in agreement with the results presented in the main text for the mixed-field TFIM model and, therefore, reflect the universality of the result. We assert that the initial (pre-scrambling) growth rate of the Krylov complexity is always increased in the presence of non-locality.
4 Conclusions
To conclude, we have explored the behavior of K-complexity in non-local spin chains. The purpose of this study is to delineate what kind of novel features this quantity can capture in the presence of non-local couplings. Over the course of this work, the K-complexity has emerged as a quantity of prime importance as it shows signs of non-locality in the system. Let us recount the major findings in the following.
4.1 Discussion of results
Lanczos sequence
The saturation and initial growth of Lanczos coefficients is greater in non-local TFIM Hamiltonian relative to the local case, which reflects the faster information scrambling in non-local systems. The local chaotic dynamics, along with long-range non-local interactions, are sufficient to give rise to fast scrambling. This is because the Lanczos sequence for chaotic Hamiltonian with local interactions saturates at a comparatively smaller value than its non-local chaotic counterpart. The large saturation value of Lanczos coefficients are the result of non-local interactions in the system, but they can equivalently be said to be the result of fast scrambling as well. The fluctuation in the non-local cases is also suppressed, which is a result of integrability breaking due to non-local interaction. As we see later, this results in a higher saturation value of Krylov complexity.
Krylov Complexity
Apart from common features of Krylov complexity, such as the transition from exponential growth to linear and the late time saturation, we find that the initial time growth rate is greater in non-local cases, showing the fast scrambling in the system. The higher growth rate in Krylov complexity for the non-local cases is a result of the initial faster growth of Lanczos coefficients in non-local cases, thus indicating faster scrambling due to the non-local interaction. Furthermore, since non-locality results in integrability breaking, we also find a large saturation value in Krylov complexity satisfying the bound in terms of the Krylov dimension. It is important to note that although the non-local interactions with the integrable local part of Hamiltonian show level statistics that of a chaotic model, it does not attain K-complexity saturation value similar to that of a chaotic case. In terms of associated Lanczos coefficients, the non-local integrable case has larger fluctuations as compared to chaotic cases.
In summary, our findings provide compelling evidence that operator growth, as characterized by K-complexity, exhibits a significantly greater magnitude in non-local systems when compared to their local counterparts. An increase in the non-locality parameter leads to a heightened scrambling rate in the system, as evident from both the Lanczos spectrum and the K-complexity profile. Furthermore, our observations indicate that the escalation of the non-locality parameter results in the breakdown of integrability and a transition towards chaotic behavior in the system. The integrability breaking due to non-local interaction is also evident from our study level statistics as well as spectral form factor (see Appendix A.1 and A.2).
4.2 Outlook
In future work, we would like to understand the relation between non-locality and fast scrambling better. This requires a careful investigation of scrambling time which is logarithmic in system size for fast scramblers, therefore requires investigation on system with large sizes. [^4] We leave this for future study due to the lack of resources to perform computation for higher dimensions. It would be interesting find the degree of non-locality require for fast scramblers by studying model with power-law decaying interactions. In ref PhysRevA.102.022402 , the authors shown that the fast scrambling is prohibited in models with a generic all-to-all term with prefactor $\sim 1/N^{\gamma}$ if $\gamma>1/2$. It would be interesting if this bound can be verified from operator growth using K-complexity.
Although all-to-all interactions are thought to be essential for fast scrambling belyansky2020minimal , certain systems demonstrate fast scrambling even in the absence of such interactions ikedaCommentsMinimalModel2021 . These models consider spin chains with next-to-nearest neighbor interactions in a specified combination of some sectors. The spread of quantum information in similar models with mild non-locality are studied in literature ecclesSpeedingSpreadQuantum2021a . These study show that although non-locality increases the rate of information spreading, but in lattice models these rates are suppressed. An extension of this work may consider K-complexity in such mild non-local models.
Acknowledgements.
We wish to thank Sumilan Banerjee, Rathindra Nath Das, Bidyut Dey, Johanna Erdmenger, Mario Flory, Aaron Friedman, Michal Heller, Chethan Krishnan, Manas Kulkarni, Subroto Mukerjee, Pratiik Nandy, and Aninda Sinha for various useful discussions and comments about this and related works. The work of AB is supported by the Polish National Science Centre (NCN) grant 2021/42/E/ST2/00234.
Appendix A Appendices: Other probes of chaos: level statistics and spectral form factor
A.1 Level Statistics
In quantum chaos, Random Matrix Theory (RMT) is a powerful tool that accurately describes the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behavior. For quantum Hamiltonians associated with integrable classical systems, the Berry-Tabor conjecture suggests that their level statistics follow a Poisson distribution 10.1007/978-3-0348-8266-8_36 .
However, for quantum Hamiltonians whose classical counterparts are chaotic, the Bohigas-Giannoni-Schmit conjecture proposes that their level statistics should fall into one of the three classical ensembles of RMT PhysRevLett.52.1 . These ensembles correspond to Hermitian random matrices with independently distributed entries as follows:
  • Gaussian Orthogonal Ensemble (GOE): Corresponding to systems with real random variables as matrix entries.
  • Gaussian Unitary Ensemble (GUE): Corresponding to systems with complex random variables as matrix entries.
  • Gaussian Symplectic Ensemble (GSE): Corresponding to systems with quaternionic random variables as matrix entries.
These ensembles provide a robust framework for understanding the statistical behavior of quantum systems with chaotic classical dynamics. Consider $e_{n}$ be an ordered set of energy levels and the nearest-neighbor spacing, $s_{n}=e_{n+1}-e_{n}$. Now, define the ratios $\tilde{r}_{n}$ as
$\tilde{r}_{n}=\frac{\text{min}(s_{n},s_{n-1})}{\text{max}(s_{n},s_{n-1})}=\text{min}\left(r_{n},\frac{1}{r_{n}}\right)$ (21)
where
$r_{n}=\frac{s_{n}}{s_{n-1}}\,.$ (22)
The distribution $P(r)$ for random matrix ensembles Wigner ensembles (GOE, GUE, and GSE) is shown to PhysRevLett.110.084101
$P(r)=\frac{1}{Z_{\beta}}\frac{(r+r^{2})^{\beta}}{(1+r+r^{2})^{1+(3/2)\beta}}$ (23)
with $Z_{\beta}$ the normalization constant. The Wigner ensembles are distinguished by their Dyson index ($\beta=1,2$ and $4$ respectively) in distribution Eq. (23). For a Poissonian distribution of level-spacings, the distribution of $r$ is given by
$P(r)=\frac{1}{(1+r)^{2}}$ (24)
The distribution $P(r)$ and $P(\tilde{r})$ are related to one another by $P(\tilde{r})=2P(r)\Theta(1-r)$. The mean value $\langle\tilde{r}\rangle$ associated with different distribution (listed in Table A.1) can be used to distinguish between integrable and chaotic systems.
Ensembles Poisson GOE GUE GSE
$\langle\tilde{r}\rangle$ $12\ln 2-$ 1 $4-2\sqrt{3}$ $2\frac{\sqrt{3}}{\pi}-\frac{1}{2}$ $\frac{32}{15}\frac{\sqrt{3}}{\pi}-\frac{1}{2}$
$\approx 0.38629$ $\approx 0.53590$ $\approx 0.60266$ $\approx 0.67617$
Figure 9: Probability distribution functions of the $\tilde{r}$ statistics for fixed $Z$-reflection symmetry blocks and Parity block of the Hamiltonian (1), computed for $N=13$ spins in the $P=+1$, and $z=+1$ sector in (a) Integrable and (b) Chaotic limits. The results are obtained from averaging over 100 ensembles with $\epsilon$ drawn from normal distribution with mean 00 and deviation $10^{-4}$. The mean level spacing ratio obtained for the integrable and chaotic limits are 0.38733 and 0.53433 respectively.
Results
In Fig. 9, we showed the probability distribution functions of the $\tilde{r}$ statistics of the local TFIM Hamiltonian (1) in integrable and chaotic regimes. We compare the results for both $P(\tilde{r})$ and $\langle\tilde{r}\rangle$ with the analytical results for Poisson and GOE symmetry class. The numerical result are in good agreement with those of analytical.
Figure 10: Probability distribution functions of the $\tilde{r}$ statistics for fixed $Z$-reflection symmetry blocks and Parity block of the non-local Hamiltonian (1), computed for $L=13$ spins in the $P=+1$, and $z=+1$ sector and non-locality parameter $\gamma=0.5$ in (a) Integrable and (b) Chaotic limits of local Hamiltonian. The results are obtained from averaging over 100 ensembles with $\epsilon$ drawn from a normal distribution with mean 00 and deviation $10^{-4}$. The mean level spacing ratio obtained for integrable and chaotic limits are 0.53784 and 0.52738 respectively.
In Fig. 10, we showed the probability distribution function of the $\tilde{r}$ statistics for non-local TFIM Hamiltonian (see Sec. 3.1) computed for $L=13$ spins. As reported in belyansky2020minimal , we find that non-locality term changes the integrable model to chaotic for $|\gamma|\gtrsim 0.25$.
Figure 11: Probability distribution functions of the $\tilde{r}$ statistics for Transverse mixed field Ising model for varying exponent $\alpha$. Refer Sec. 3.2 for parameter values.
In Fig. 11, we shown level statistics distribution for Transverse mixed field Ising model discussed in Sec 3.2 for varying non-local exponent $\alpha$. It is clear from here that it is hard to distinguish these different degrees of non-locality from the level statistics, while the initial growth rate of K-complexity can quite clearly distinguish them. The reason is that in K-complexity, we can distinguish between different degrees of non-locality during the pre-scrambling regime. At the same time, the level statistics can capture only the details about the nature of the spectrum, and there is no way to make such a timewise distinction.
A.2 Spectral Form Factor
To compute the spectral form factor, we use the analytically continued partition function
$Z(\beta,t)\equiv\text{Tr}\left(e^{-\beta H-iHt}\right)$ (25)
of the usual partition function $Z(\beta)=\text{Tr}\left(e^{-\beta H}\right)$. The function $Z(\beta,t)$ is an erratic function of time and, therefore oscillates largely at late times. The fluctuations are studied by considering the normalized squared quantity, the so-called spectral form factor (SFF) cotlerBlackHolesRandom2017 ; PhysRevD.95.126008 ; winerSpectralFormFactor2022
$g(\beta,t)=\left|\frac{Z(\beta,t)}{Z(\beta)}\right|^{2}=\frac{1}{Z(\beta)^{2}}\sum_{m,n}e^{-\beta(E_{m}+E_{n})}e^{i(E_{m}-E_{n})t}$ (26)
where $E_{n}$ is eigenvalue of energy eigenstate $|n\rangle.$ The long-time average can be written as
$\lim_{T\rightarrow\infty}\frac{1}{T}dt\left|\frac{Z(\beta,t)}{Z(\beta)}\right|^{2}=\frac{Z(2\beta)}{Z(\beta)^{2}}$ (27)
where we assumed the system Hamiltonian $H$ to be non-degenerate. Due to the erratic nature of $Z(\beta,t)$, usually the ensemble average over large Hamiltonians is considered. We define disorder-averaged analogs of SFF in Eq. (26) as
$g_{\text{ann}}(\beta,t)$ $=\frac{\mathbb{E}[|Z(\beta,t)|^{2}]}{\mathbb{E}[|Z(\beta,0)|^{2}]}$ (28)
$g_{\text{que}}(\beta,t)$ $=\mathbb{E}\left[\left|\frac{Z(\beta,t)}{Z(\beta,0)}\right|^{2}\right]$ (29)
which are called annealed SFF and quenched SFF, respectively. Throughout this paper, we use the notatin $\mathbb{E}[.]$ to denote the ensemble average. In this work, we will be working with annealed SFF along with $\beta=0$, meaning that we are taking the disorder average separately in the numerator and denominator. We will further call $g_{\text{ann}}(\beta,0)=g(t)$, sometime this itself is referred to as the spectral form factor, which has information about the correlations of eigenvalues at different energy separations. To understand the time profile of SFF, we split the $|\mathbb{E}[|Z(\beta,t)|^{2}]$ into two parts as
$|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}+\left(\mathbb{E}\left[|\text{Tr}(e^{-iH(\beta+it)})|^{2}\right]-|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}\right)$ (30)
The first term is a disconnected part of the SFF, which comes solely from the average density of states. The second part is a connected part which contains the information on the correlation between energy levels. According to random matrix universality, an ensemble of quantum chaotic Hamilontians will generically have the same connected SFF as the canonical Gaussian ensembles of RMT. The conjectured universal profile of the SFF of the GUE contains three distinct regimes winerSpectralFormFactor2022 ; Cipolloni_2023 ; brezinSpectralFormFactor1997 :
  • Initially the value drops quickly, through a region we call the slope, to a minimum, which is called the dip. The early time dip in profile comes from the disconnected part of the SFF therefore, its precise shape is non-universal. It is the result of loss of constructive interference in different terms of $\text{Tr}e^{-iH(\beta+it)}$ which acquire different phase factors as $t$ increases.
  • After a dip, the values increase roughly linearly which is what we call the ramp. The ramp is due to the repulsion between eigenvalues that are far apart in the spectrum, which is also well-known in quantum chaotic systems. Therefore a linear ramp is often taken as a defining signature of quantum chaos.
  • The plateau, which occurs at late times, results from the discreteness of the spectrum.
Figure 12: A log-log plot of TFIM Hamiltonian SFF $g(t;\beta=0)$, plotted against time for $L=11$ in (a) Integrable (b) Chaotic limits. The value at late times, which is equal to plateau height $g_{p}$, matches with $Z(2\beta)/Z(\beta)^{2}$ (shown in black dashed line). The average is taken over $50,000$ samples drawn from the random value of $\epsilon$ taken from a normal distribution with mean $\mu=0$ and standard deviation $\sigma=0.01$ in case of chaotic limit. The dash-dot line shows the linear fit in the ramp region, showing the linear increase.
The overall profile for GOE and GSE symmetry class also follows a slope-dip-ramp-plateau pattern apart from a few details, such as the shape of the ramp and the plateau. For example, the SFF of the GUE symmetry class has a sharp kink, while in the GOE symmetry class, the kink is smoothed Cipolloni_2023 .
Figure 13: A log-log plot of non-local TFIM Hamiltonian SFF $g(t;\beta=0)$, plotted against time for $L=11$, and $\gamma=0.5$ in (a) Integrable (b) Chaotic limits of local TFIM Hamiltonian. The value at late times, which is equal to plateau height $g_{p}$, matches with $Z(2\beta)/Z(\beta)^{2}$ (shown in black dashed line). The average is taken over $50,000$ samples drawn from the random value of $\epsilon$ taken from normal distribution with mean $\mu=0$ and standard deviation $\sigma=0.01$. The dash-dot line shows the linear fit in the ramp region, showing the linear increase.
Results
[…]
In Fig. 13, we show the SFF of non-local TFIM Hamiltonian for $L=11$, and $\gamma=0.5$ in the Integrable and Chaotic limits. We took the disordered average over 50,000 samples drawn from the random value of $\epsilon$. We find slope-dip-ramp-plateau patterns in both limits, integrable and chaotic, indicating that the non local interaction makes the Hamiltonian chaotic even when its local counterpart is in the integrable regime. It is also worth noting that there is no clear way to distinguish between the two non-local plots here. On the other hand, as described in the main text, by looking at the saturation values of the non-local K-complexity plots, we can still distinguish between which was integrable and which was not in the local version.